OpenStudy (anonymous):

Renna pushed the button for the elevator to go up, but it would not move. The weight limit for the elevator is 450 kilograms, but the current group of passengers weighs a total of 750 kilograms. Renna wants to determine how many 70-kilogram passengers need to get off the elevator. Let p represent the number of excess passengers. Write an inequality to determine the number of passengers who need to get off the elevator to meet the weight requirement.

OpenStudy (perl):

You can use the inequality total_weight - 70 * (# excess passengers) <= weight_limit

OpenStudy (anonymous):

Here's my solution: let limit weight be q= 450 let passengers weight be p=750 Therefore, the passengers are 300 over the limit this is z p-z is equal to q We know that there are 10.7 approx. 11 passengers on the elevator that make it 750 by 750/70 Therefore, 300/70 gives use how many passengers should leave 4.2857 approx. 5. So 11-5= 6 passengers should stay. This all can be represented by p - ((p-q)/70) which should be less than or equal too q. This depends if you round or not. Message me for more info.

OpenStudy (perl):

i think its simpler to use the inequality 750 - 70*p <= 450 assuming that p >0